Turkey Company is seeking to borrow some money on a long-term basis. This firm is considering pursuing one of two alternatives: (1) issuing fifty bonds with a face amount of $1,000 each and (2) signing a long-term note payable for $55,000. Assume that both alternatives would provide cash to Turkey on December 31, Year 1.

The terms of the two alternatives are presented below:

  1. The bonds would have a coupon rate of 9% and a market rate [effective rate] of 8%. The bonds would pay interest semiannually on June 30 and December 31.   The bonds would have a maturity date of December 31, Year 11.
  2. The note payable would have an interest rate of 8% and would be paid off in 24 equal semi-annual installments starting six months after December 31, Year 1.

Required:

  1. How much cash would each alternative provide to Turkey on Dec. 31, Year 1?

Bonds $_________

Note $__________

  1. What interest expense would the firm recognize in Year 2 under both alternatives?

Bonds $_________

Note $__________

  1. How much cash would Turkey have to pay out during Year 3 if it chose to issue the Bond Payable?

­­­­­$____________

  1. How much cash would Turkey have to pay out in Year 3 if it chose to borrow the money using the long-term Note Payable?

$____________

  1. What wold the carrying value of the Note Payable be on December 31, Year 5, after the eighth loan payment that would be made on that date?

$________________
Answer #1
Cash proceeds-Bonds:
Cash proceeds of the bond=Present value of interest payment+Present value of principal repayment
Interest payment=Bond face value*Coupon rate=2090000*11%=$ 229900
Bond face value=Number of bonds*Face amount=50*1000=$ 50000
Interest payment per semi-annual period=50000*9%*(6/12)=50000*4.5%=$ 2250
Issued on Dec 31, Year 1
Maturity date on Dec 31,Year 11
Life of the bond=10 years
Number of semi-annual periods=10*2=20 years
Discount rate=Market rate for 6 months=8%*(6/12)=4%
Present value of interest payments=94050* PVIAF @ 4% for 20 years=2250*13.59033=$ 30578
Present value of principal repayment at the end of 20th year=50000*0.45639=$ 22820
Cash proceeds of the bond=30578+22820=$ 53398
Cash proceeds-Note:
Cash proceeds=Face value of note=$ 55000
Bond amortization schedule:
Date Cash
paid
Interest
expense
Premium
amortizaton
Carrying
amount
of bonds
(Interest payment per semi-annual period) (Previous
carrying
amount)*4%
(Previous
carrying
amount-
c)
a b c=a-b
Dec 31,Year 1 53398
June 30,Year 2 2250 2136 114 53284
Dec 31,Year 2 2250 2131 119 53165
June 30,Year 3 2250 2127 123 53042
Dec 31,Year 3 2250 2122 128 52914
June 30,Year 4 2250 2117 133 52780
Dec 31,Year 4 2250 2111 139 52641
June 30,Year 5 2250 2106 144 52497
Dec 31,Year 5 2250 2100 150 52347
June 30,Year 6 2250 2094 156 52191
Dec 31,Year 6 2250 2088 162 52028
June 30,Year 7 2250 2081 169 51859
Dec 31,Year 7 2250 2074 176 51684
June 30,Year 8 2250 2067 183 51501
Dec 31,Year 8 2250 2060 190 51311
June 30,Year 9 2250 2052 198 51114
Dec 31,Year 9 2250 2045 205 50908
June 30,Year 10 2250 2036 214 50695
Dec 31,Year 10 2250 2028 222 50472
June 30,Year 11 2250 2019 231 50241
Dec 31,Year 11 2250 2009 241 50000
Discount rate for note=8%*(6/12)=4% (Since equal semi-annual installment)
Monthly installment payment of note=Face value of note/Discount factor at 4% for 24 years=55000/15.24696=$ 3607.8=$ 3608
Note amortization schedule:
Date Installment paid Interest
expense
Principal repayment Carrying
amount
of note
(Previous
carrying
amount)*4%
(Previous
carriying
amount-
c)
a b c=a-b
Dec 31,Year 1 55000
June 30,Year 2 3608 2200 1408 53592
Dec 31,Year 2 3608 2144 1464 52128
June 30,Year 3 3608 2085 1523 50605
Dec 31,Year 3 3608 2024 1584 49021
June 30,Year 4 3608 1961 1647 47374
Dec 31,Year 4 3608 1895 1713 45661
June 30,Year 5 3608 1826 1782 43879
Dec 31,Year 5 3608 1755 1853 42026
June 30,Year 6 3608 1681 1927 40099
Dec 31,Year 6 3608 1604 2004 38095
June 30,Year 7 3608 1524 2084 36011
Dec 31,Year 7 3608 1440 2168 33844
June 30,Year 8 3608 1354 2254 31589
Dec 31,Year 8 3608 1264 2344 29245
June 30,Year 9 3608 1170 2438 26807
Dec 31,Year 9 3608 1072 2536 24271
June 30,Year 10 3608 971 2637 21634
Dec 31,Year 10 3608 865 2743 18891
June 30,Year 11 3608 756 2852 16039
Dec 31,Year 11 3608 642 2966 13072
June 30,Year 12 3608 523 3085 9987
Dec 31,Year 12 3608 399 3209 6779
June 30,Year 13 3608 271 3337 3442
Dec 31,Year 13 3608 166 3442 0
Interest expense for year 2: (Refer amortization table)
Bond=2131+2127=$ 4258
Note=2200+2144=$ 4344
Cash to be paid in year 3: (Refer amortization table)
Bond=2250+2250=$ 4500
Note=3608+3608=$ 7216
Carrying value of note as on Dec 31,Year 5=$ 42026